# Diffence: Fencing Membership Privacy With Diffusion Models

### v1

@@ 4.2 Experimental Results
Figure 6: Attack accuracy and AUC on three datasets against DenseNet under Scenario 1. We report the highest attack accuracy and attack AUC across all attacks. The prediction accuracy delta indicates the prediction accuracy gap compared to the undefended models, with negative numbers indicating a decrease in model utility.
TABLE III: Average attack accuracy and AUC across three datasets under three scenarios on ResNet. The optimal defense results in each case are highlighted in bold.
Defenses
Prediction Accuracy Delta (%)
w/o Diffence w/ Diffence (Scenario 1) w/ Diffence (Scenario 2) w/ Diffence (Scenario 3)
Attack AUC (%)
Attack Accuracy (%)
Attack AUC (%)
Attack Accuracy (%)
Attack AUC (%)
Attack Accuracy (%)
Attack AUC (%)
Attack Accuracy (%)
Undefended
0
79.57
77.22
71.33 67.63
73.17
69.9
72.86
69.98
@@
-1.89
61.03
58.92
58.53 56.38
59.12
57.59
58.99
57.51
AdvReg
-6.63
61.77
59.54
59.49 57.41
59.57
57.42
59.53
57.43
@@
-0.27
80.08
76.67
72.63 67.85
73.98
69.44
73.63
69.46
RelaxLoss
0.11
75.29
72.29
69.06 65.35
69.73
67.28
69.17
67.20
DP-SGD
-9.10
56.44
55.75
55.48
54.8
55.47 54.46
55.33
54.56
Memguard
0
72.50
68.49
67.41 64.63
68.90
65.81
69.10
65.81
4.2.1 Comparison to Baselines
We conducted an extensive evaluation of our method across three datasets. For each dataset, we evaluated the performance of models across seven different cases including one without any protection and six others, each employing a different defense mechanism. We tested the impact of our method on the models’ test accuracy and the changes in attack accuracy and attack AUC under three scenarios introduced in Section 4. Our results show that our method consistently enhanced privacy protection across all settings without compromising the models’ utility. We discuss the specific effects of our method below.
Figure 5 and Figure 6 shows the performance of our method against ResNet and DenseNet under Scenario 1, respectively. We can observe that our approach consistently reduces the attack’s AUC and accuracy across all cases, particularly against those methods that preserve model usability but offer only limited protection. For instance, when applied to ResNet, our method managed to decrease the attack accuracy for undefended, HAMP, RelaxLoss, and Selena models by 12.4%, 11.2%, 8.9%, and 5.7%, respectively on average across three datasets. Similarly, the attack AUCs showed reductions of 10.1%, 9.1%, 7.9%, and 3.9% after employing our approach. The experimental results demonstrate that our method can significantly reduce privacy threats, even with reliance on only a small amount of additional data (1000 non-members in our experiments). Our experiments suggest that the optimal defense practice is combining our method with recent utility-preserving defenses. This combination leverages their advantages in maintaining utility while enhancing their privacy protection. For instance, the integrations of Diffence with MemGuard, Selena, and RelaxLoss respectively achieved the best privacy-utility trade-offs on three different datasets when applied to DenseNet.
Table III presents the experimental results using ResNet under all three different scenarios. We show the average attack AUC and accuracy reduction of our method across three datasets in each scenario. The average defense effectiveness in Scenarios 2 and 3 is slightly weaker than in Scenario 1, due to stronger assumptions placed on the defender in these scenarios. However, it is observable that they still achieve significant privacy enhancements across all settings without compromising the model’s utility. In the research of MIA defense, balancing sufficient privacy protection with model utility has always been one of the most important objectives. Our results suggest that our method represents a significant step forward towards the ideal defense.
4.2.2 Effectiveness on Attack TPR and TNR Metrics
Figure 7: Attack TPR and FPR on SVHN dataset We report the highest attack TPR and attack FPR across all attacks.
Following the practice of Carlini et al. [29], we measure the reliability with which an adversary can violate the privacy of even a few users in a sensitive dataset using metrics such as TPR at 0.1% FPR and TNR at 0.1% FNR. We report on the defensive effectiveness of our method under the third scenario against these metrics on SVHN, employing six attacks including LiRA. As shown in Figure 7, our method significantly reduces the attack TPR and TNR in most cases. For instance, our method reduced the attack TPR against the undefended model and HAMP by 84.7% and 61%, respectively. Our approach has a relatively smaller impact on methods that already exhibit low TPR, but these methods often compromise the model’s utility. Therefore, our method can be advantageously combined with privacy-preserving defenses.
4.2.3 Improving Performance Beyond Previous State-of-the-Art
TABLE IV: Performance of our approach on CIFAR-100. Here we directly adopted settings from previous papers.
SELENA
defenses
Training accuracy (%)
Test accuracy (%)
Attack AUC (%)
Attack accuracy (%)
Undefended / +ours
99.98 / 99.98
78.19 / 78.19
77.72 / 63.4
74.02 / 60.73
SELENA / +ours
78.19 / 78.19
74.47 / 74.47
56.56 / 53.37
55.18 / 53.08
RelaxLoss
defenses
Training accuracy
Test accuracy
Attack AUC
Attack accuracy
Undefended / +ours
65.67 / 65.67
32.99 / 32.99
71.69 / 59.5
67.58 / 58.04
RelaxLoss / +ours
52.68 / 52.68
35.7 / 35.7
59.93 / 54.95
57.43 / 53.61
HAMP
defenses
Training accuracy
Test accuracy
Attack AUC
Attack accuracy
Undefended / +ours
92.98 / 92.98
58.51 / 58.51
68.47 / 67.4
67.09 / 66.02
HAMP / +ours
83.38 / 57.25
83.38 / 57.25
65.72 / 62.69
63.51 / 60.13
For a fair and accurate comparison, we adopted the experimental settings deployed in their original papers to assess the effectiveness of our proposed method. The results are shown in Table IV. Note that our approach maintains the original predicted labels of the model. While this does not eliminate the privacy risks posed by the accuracy gap between training and test sets, it significantly mitigates the risk of privacy leakage arising from other inconsistent prediction behaviors.
Consider the instance of SELENA: the undefended model exhibits a 21.79% gap (99.98% - 78.19%) between training and test set accuracies. Consequently, a simplistic gap attack — assuming all correctly classified samples as members — could attain an attack accuracy of 60.9%. In contrast, our defense reduces the best performance of other attacks to an accuracy of just 60.73%. This implies that attackers can gain no improvement over the baseline attack, a substantial reduction from the 4.7% improvement possible with SELENA.
Beyond merely lowering attack accuracy, our method can combined with other defenses, achieving reduced attack AUC and accuracy across all scenarios without compromising the model’s accuracy. Our experiments demonstrate that our method can be applied to the optimal defense models proposed in prior works, enhancing their privacy protection.
4.2.4 Defend with Stronger Diffusion Models
In our previous experiments, we assumed that the training set size of the defender’s diffusion model and the size of the classifier’s training set are the same. Furthermore, to ensure no decrease in utility, we only select reconstructed samples that match the original sample labels.
Interestingly, however, we discovered that when a stronger diffusion model is employed in the defense—specifically, a model trained on a dataset larger than that of the classifier, directly averaging the predictions of all generated samples can enhance both the classifier’s privacy and its utility.
We utilized a publicly available diffusion model trained on 50,000 samples from CIFAR-10 to defend our classifiers, which were trained on 25,000 samples. The results, as shown in the Table V, indicate that in all cases, the classifiers not only achieved enhanced privacy but also exhibited higher test accuracy. Given that diffusion model training does not require labeled data, it can be performed on a large, public, unlabeled dataset. We leave a more detailed discussion on how diffusion models can improve classifiers’ utility for future work.
TABLE V: Performance of our approach on CIFAR-10 when using a stronger diffusion model.
Defenses
Test Accuracy (%)
Attack Accuracy (%)
Attack AUC (%)
Undefended / +Ours
81.4 / 83.18
72.98 / 66.19
75.85 / 69.3
AdvReg / +Ours
78.4 / 80.32
60.31 / 57.52
60.99 / 58.17
DPSGD / +Ours
74.66 / 75.72
55.31 / 54.38
55.3 / 54.47
SELENA / +Ours
79.2 / 79.8
60.36 / 57.17
60.08 / 57.02
RelaxLoss / +Ours
81.24 / 82.08
65.13 / 61.48
69.52 / 63.7
HAMP / +Ours
81.12 / 82.74
65.2 / 62.5
69.44 / 62.16
4.3 Ablation study
As introduced in Section 3, our approach employs a diffusion model to reconstruct samples, protecting the membership privacy of the original samples. We generate multiple samples and, based on the defense objectives and the information available to the defender, select the prediction of the most suitable sample as the final output. This process involves considering several factors, including the choice of interval, the number of reconstructions $N$ generated for each sample, and the diffusion step $T$. In this section, we delve further into discussing the impact of the parameters and sample selection strategy.
4.3.1 Sample Selection Strategies
Figure 8: Cumulative Distribution Function (CDF) of the number of generations required for reconstructions to fall within different intervals when $T=60$ on CIFAR100. The ’percentile’ indicates the anticipated probability of reconstructions falling within the given interval, as determined by analyzing the samples available to the defender.
In Scenario 1, the main idea of our method involves setting an interval within which the prediction logits of both members and non-members are encouraged to fall. This strategy aims to align the prediction distributions of members and non-members towards a range in the middle. It is important to note that our method uses a fixed number of reconstructions $N$ and diffusion steps $T$ as the basis for selecting the optimal interval. Given the fixed $N$, not all samples can fall within the chosen interval. For these outliers, our approach is to select the closest sample.
@@ 4.3.1 Sample Selection Strategies
Figure 9 depicts attack AUC and accuracy on three datasets in relation to the different JS divergence between member and non-member prediction distributions when our method is solely applied. Each point in the figure corresponds to a selectable interval, where choosing different intervals results in varying JS divergences. A key observation is the negative correlation between JS divergence and privacy protection – smaller JS divergences typically indicate stronger membership privacy protection. These results substantiate the effectiveness of our interval selection strategy.
(a) CIFAR10
(b) CIFAR100
(c) SVHN
Figure 9: Attack AUC and accuracy under different levels of Jensen-Shannon (JS) divergence between member and non-member prediction distributions. We tested it on three dataset using ResNet and set diffusion steps $T=40$ and the number of reconstructions $N$ to 50.
4.3.2 Effect of Hyperparameters
@@
Figure 10: Attack AUC and Attack Accuracy Against Our Defended Models with Different Numbers of Generated Samples $N$ and Diffusion Steps $T$. We evaluate using ResNet on CIFAR100 dataset, varying $N$ from 25 to 200 and $T$ from 10 to 80.
We then varied the values of $N$ and $T$ to observe the changes in defense performance. Note that our method does not alter the prediction labels, and therefore, does not impact the model’s test accuracy.
@@
5 Discussion
5.1 Overhead of Different Defenses
TABLE VI: Training and inference overhead comparison of different defenses.For our method, parameters are set to their default values used in our experiments with $N=50$ and $T=40$.
Defenses
CIFAR10 CIFAR100 SVHN
Training overhead
Inference overhead
Training overhead
Inference overhead
Training overhead
Inference overhead
Undefended
0.6h
25.8ms
0.6h
24ms
0.4h
24ms
SELENA
10.9h
25.8ms
11.9h
24ms
9.3h
24ms
AdvReg
10.9h
25.8ms
12.6h
24ms
6.8h
24ms
Hamp
0.8h
25.8ms
0.77h
24ms
0.5h
24ms
RelaxLoss
0.53h
25.8ms
0.6h
24ms
0.44h
24ms
DP-SGD
0.83h
25.8ms
0.94h
24ms
0.56h
24ms
Memguard
732ms
612ms
586ms
Ours
245ms
221ms
241ms
We compared the overhead of 7 defense methods. Among them, Hamp, SELENA, RelaxLoss, AdvReg, and DP-SGD are deployed in the training phase, and memguard and our method are deployed in the inference phase. All models are trained with 100 epochs. Times are measured on a single NVIDIA RTX-8000 GPU with 40GB memory.
As shown in Table VI, the training time of the defense deployed in the training phase is 1x $\sim$ 21x relative to that of the undefended model. Memguard and our method have no impact on the training time. Their overhead is at inference time. Among them, our method outperforms Memguard in terms of running time in the inference phase, as Memguard needs to solve complex optimization problems for each sample.
Note that We used the standard DDIM sampling method in our approach. With the development of fast sampling technology, the overhead of our defense might be further reduced. We leave the use of fast sampling in our defense framework to future work.
6 Limitation
While our method demonstrates significant efficacy in enhancing privacy protection, it is important to acknowledge certain limitations inherent in our approach. A key aspect of our approach is the use of diffusion models. In many cases, off-the-shelf diffusion models are sufficient and can be directly employed without additional training. This availability significantly reduces the resources and time required, making our method more accessible and practical. However, in scenarios where the dataset requiring privacy protection does not have a readily available corresponding diffusion model, training or fine-tuning becomes necessary.
It is essential to note that once fine-tuned, the diffusion model is not limited to a singular application but can be leveraged for various other tasks and scenarios. This versatility offers a significant advantage, as the investment in training the model can be amortized over multiple uses. For instance, a fine-tuned diffusion model on a medical dataset could be utilized not only for privacy protection but also for tasks such as data augmentation, anomaly detection, or even generating synthetic data for research purposes [34, 35].
Nevertheless, the initial requirement of model training or fine-tuning, particularly on private or sensitive datasets, remains a critical factor to consider. This process involves careful handling and management of the data to ensure that the privacy of the data subjects is not compromised during the model development phase.
In summary, while our method requires an initial investment in training or fine-tuning the diffusion model, the subsequent benefits and applications of the model extend well beyond the scope of privacy protection, offering a range of possibilities in data utilization.
7 Related Work
@@
7.1 Membership Inference Attacks
Shokri et al. [4] introduced a black-box MIA that employs a shadow training technique to train an attack model to differentiate the model’s output, categorizing it as either a member or non-member. In a different approach, Salem et al. [15] streamlined the process by training only a single shadow model, assuming the attacker lacks access to similar distribution data as the training dataset, yet still achieving notable effectiveness. Expanding on these concepts, Nasr et al. [36] presented a white-box MIA targeting ML models. For each data sample, they computed the corresponding gradients over the parameters of the white-box target classifier, utilized as features of the data sample for the purpose of membership inference.
@@ 8 Conclusion
In this study, we introduced a novel privacy defense method that addresses the challenge of membership inference attacks (MIAs) in machine learning (ML). Our approach effectively diminishes the distinction in prediction behaviors between members and non-members through input reconstruction using diffusion models. By generating multiple reconstructions and selectively utilizing predictions based on defined criteria, our method significantly narrows the prediction distribution gaps that are often exploited in MIAs.
@@
We categorized defenses into three deployment phases and, for the first time, proposed integrating different defenses to enhance overall protection. By combining our method with others, we are able to effectively address both the train-to-test accuracy gap and the prediction distribution gap. Our experiments across various datasets demonstrated the efficacy of our approach in reducing membership privacy leakage, without sacrificing the model’s accuracy.
Our method demonstrates considerable flexibility and effectiveness. However, we recognize its limitations in scenarios where specific diffusion models are not readily available, necessitating additional training or fine-tuning. Despite this, the potential for utilizing fine-tuned diffusion models in various contexts offers a substantial advantage.
In conclusion, our research contributes to the growing body of knowledge in privacy protection in ML, offering a robust, flexible solution to counter MIAs. Future work could further explore optimizing the selection process and investigating its effectiveness against a wider range of privacy threats in ML.

### v2

@@ IV-B4 Improving Performance Beyond Previous State-of-the-Art
Figure 8: Attack accuracy and AUC on two high-resolution datasets against ResNet18, where Diffence employs a publicly released diffusion model trained on ImageNet.
For a fair and accurate comparison, we adopted the experimental settings deployed in their original papers to assess the effectiveness of Diffence. In our evaluations, Diffence was configured with $N=10$ and $T=50$ in Scenario 3. The results are shown in Table IV. Note that Diffence maintains the original predicted labels of the model. While this does not eliminate the privacy risks posed by the accuracy gap between training and test sets, it significantly mitigates the risk of privacy leakage arising from other inconsistent prediction behaviors.
We observed that Diffence can effectively enhance membership privacy in most settings. For instance, for HAMP, Diffence successfully reduced attack AUC and accuracy from 67.4% and 65.9% to 64.7% and 61.5%, respectively. Diffence did not show improvement in the RelaxLoss setting, as there was no room for privacy enhancement by reducing the prediction distribution gap. In this case, a simple attack assuming all correctly classified samples as members achieved the highest attack accuracy. Our experiments demonstrate that Diffence can be applied to the best defense models proposed in prior works, further enhancing privacy by reducing the prediction distribution gap.
TABLE IV: Performance of Diffence on CIFAR-100. Here we directly adopted settings from previous papers.
Defenses Training Accuracy (%) Test Accuracy (%) Attack AUC (%) Attack Accuracy (%)
@@
Undefended / +Diffence 100.0 / 100.0 78.4 / 78.4 76.2 / 64.8 74.5 / 63.0
SELENA / +Diffence 77.3 / 77.3 75.0 / 75.0 55.5 / 51.8 55.1 / 52.7
RelaxLoss
Undefended / +Diffence 67.3 / 67.3 34.1 / 34.1 72.4 / 69.5 67.9 / 67.0
RelaxLoss / +Diffence 52.9 / 52.9 36.6 / 36.6 59.5 / 59.7 57.6 / 57.7
@@
Undefended / +Diffence 91.4 / 91.4 58.7 / 58.7 70.8 / 67.9 67.5 / 66.1
HAMP / +Diffence 78.6 / 78.6 55.6 / 55.6 67.4 / 64.7 65.9 / 61.5
IV-B5 Evaluating Diffence with Publicly Available Diffusion Models
Although Diffence requires deploying a diffusion model, the model does not need to be trained on the same dataset as the target model. Therefore, defenders can leverage off-the-shelf diffusion models, including publicly available models trained on different datasets.
To demonstrate this, we evaluated Diffence on two high-resolution datasets: CelebA [25] and UTKFace [26] in Scenario 3 using a publicly available diffusion model trained on ImageNet by Dhariwal et al. [14] with settings of $N=10$ and $T=200$. To the best of our knowledge, we are the first to conduct comprehensive MIA tests on these datasets using the original resolution without resizing, employing five attacks and seven defenses. The experimental results are shown in Figure 8. It can be observed that Diffence effectively enhances membership privacy in all cases without compromising model accuracy. On CelebA, Diffence decreases attack AUC by an average of 6.2% and attack accuracy by 6.0%. On UTKFace, Diffence decreases attack AUC by 9.5% and attack accuracy by 9.9% on average. On CelebA and UTKFace, the combinations of HAMP + Diffence and SELENA + Diffence achieved the best utility-privacy trade-offs, respectively.
We also conducted experiments on three benchmark datasets—CIFAR-10, CIFAR-100, and SVHN-using the same public diffusion model. The results demonstrate that Diffence with the pretrained diffusion model provides effective defense across all three datasets. On average, Diffence reduces the attack AUC by 4.8% and attack accuracy by 5.3%. For more details, please refer to Figure 13 in the Appendix. Additionally, the reconstructions generated by the diffusion model are nearly identical to the original samples, with only imperceptible visual differences, as shown in Figure 14.
(a) CIFAR-10
(b) CIFAR-100
(c) SVHN
Figure 9: Attack AUC and accuracy under different levels of Jensen-Shannon (JS) divergence between member and non-member prediction distributions. We tested it on three dataset using ResNet18 and set diffusion steps $T=40$ and the number of reconstructions $N$ to 50.
Figure 10: Cumulative Distribution Function (CDF) of the number of generations required for reconstructions to fall within different intervals when $T=100$ on CIFAR-100. The ’percentile’ indicates the anticipated probability of reconstructions falling within the given interval, as determined by analyzing the samples available to the defender.
IV-C Ablation study
As introduced in Section III, Diffence employs a diffusion model to reconstruct samples, protecting the membership privacy of the original samples. We generate multiple samples and select the prediction of the most suitable sample as the final output based on the defense objectives and the information available to the defender. This process involves considering several factors, including the choice of interval, the number of reconstructions $N$ generated for each sample, and the diffusion step $T$. In this section, we delve further into discussing the impact of the parameters and sample selection strategy.
IV-C1 Sample Selection Strategies
In Scenario 1, the main idea of Diffence involves setting an interval within which the prediction logits of both members and non-members are encouraged to fall. This strategy aims to align the prediction distributions of members and non-members towards a range in the middle. It is important to note that Diffence uses a fixed number of reconstructions $N$ and diffusion steps $T$ as the basis for selecting the optimal interval. Given the fixed $N$, not all samples can fall within the chosen interval. For these outliers, we select the sample that is closest to the interval.
@@ IV-C1 Sample Selection Strategies
Figure 9 depicts attack AUC and accuracy on three datasets in relation to the different JS divergence between member and non-member prediction distributions when Diffence is solely applied. Each point in the figure corresponds to a selectable interval, where choosing different intervals results in varying JS divergences. A key observation is the negative correlation between JS divergence and privacy protection – smaller JS divergences typically indicate stronger membership privacy protection. These results substantiate the effectiveness of our interval selection strategy.
IV-C2 Effect of Hyperparameters
@@
Figure 11: Attack AUC and attack accuracy against Diffence with different numbers of generated samples $N$ and Diffusion Steps $T$ on average across three datasets. We varies $N$ from 2 to 50 and $T$ from 40 to 260.
We then varied the values of $N$ and $T$ to observe the changes in defense performance. Note that Diffence does not alter the prediction labels, and therefore, does not impact the model’s test accuracy.
@@
V Discussions
V-A Overhead of Different Defenses
Our experiments include seven defense methods, where HAMP, SELENA, RelaxLoss, AdvReg, and DPSGD are training-phase defenses, and MemGuard and Diffence operate during the inference phase. We evaluated the overhead of these defenses during training and testing based on their deployment phase. The training and inference overhead were measured on a single NVIDIA RTX-8000 GPU with 40GB of memory. Here, we primarily discuss the inference overhead, with Appendix A-D reporting the results on training overhead comparison.
TABLE V: Inference overhead comparison of different defenses.
Defenses CIFAR-10 (ms) CIFAR-100 (ms) SVHN (ms)
Undefended 26 24 24
Memguard 732 612 586
Diffence 82 82 81
MemGuard generates adversarial noise for each sample’s confidence vector to deceive the attacker’s model, necessitating the solving of complex optimization problems for each sample. In contrast, Diffence, which involves calculating optimal intervals for samples before inference, is a one-time task that does not add overhead during the inference phase. The primary overhead of Diffence comes from processing the original samples with a diffusion model. To enhance this process, we utilize Denoising Diffusion Implicit Models (DDIM) [45], which streamline the diffusion process by providing a means to skip sampling steps and directly estimate the denoised image from noisy observations, significantly reducing computational demands. Table V reports the inference times for the undefended model, MemGuard, and Diffence on three benchmark datasets. We measured the inference overhead by performing inference on 100 random member and non-member samples and taking the average. As shown in Table V, Diffence outperforms MemGuard in terms of inference phase running time and only adds an average of 57ms to the inference time compared to the undefended model, highlighting Diffence’s efficiency and low overhead upon deployment.
V-B Pre-training MIA Defenses
In Section I, we categorized MIA defenses into three classes based on where they kick in during the ML pipeline: training phase defenses, pre-inference defenses, and post-inference defenses. One could also consider techniques that work during ML pre-training as another class of MIA defense, e.g., techniques like dataset condensation (DC) [46]. However, we argue that these techniques tend to mitigate MIA more as a side effect than as a primary focus. Furthermore, their effectiveness in defending against MIA has been questioned in recent work [47].
Also, one could consider these pre-training techniques as a sub-class of our training phase defenses. In fact, some training stage defenses, such as SELENA and HAMP, already incorporate specifically designed dataset processing. Therefore, we did not make pre-training a separate class of MIA defenses.
VI Related Works
@@
VI-A Membership Inference Attacks
Shokri et al. [10] introduced a black-box MIA that employs a shadow training technique to train an attack model to differentiate the model’s output, categorizing it as either a member or non-member. In a different approach, Salem et al. [17] streamlined the process by training only a single shadow model, assuming the attacker lacks access to similar distribution data as the training dataset, yet still achieving notable effectiveness. Expanding on these concepts, Nasr et al. [48] presented a white-box MIA targeting ML models. For each data sample, they computed the corresponding gradients over the parameters of the white-box target classifier, utilized as features of the data sample for membership inference.
@@ VII Conclusion
In this work, we introduced Diffence, a novel defense method that addresses the challenge of membership inference attacks (MIAs) in machine learning (ML). Diffence effectively diminishes the distinction in prediction behaviors between members and non-members through input reconstruction using diffusion models. By generating multiple reconstructions and selectively utilizing predictions based on defined criteria, Diffence significantly narrows the prediction distribution gaps exploited in MIAs.
@@
We categorized defenses into three deployment phases and, for the first time, proposed integrating different defenses to enhance overall protection. Diffence can be cascaded in a plug-and-play manner with other defenses. Although our strategy requires the use of a diffusion model, it does not necessarily require the diffusion model to be specifically trained for defense; instead, off-the-shelf diffusion models can be used—even those trained on datasets different from the target model’s training dataset—thereby demonstrating considerable flexibility and effectiveness. By combining Diffence with others, we effectively address both the train-to-test accuracy gap and the prediction distribution gap. Our extensive experiments across multiple datasets have validated Diffence ’s efficacy in enhancing membership privacy without sacrificing model utility—preserving both accuracy and the meaningfulness of confidence vectors.
Acknowledgements
The work was supported in part by the NSF grant 2131910, and by DARPA under Grant DARPA-RA-21-03-09-YFA9-FP-003.
Appendix A Appendix
(a) Attack Accuracy
(b) Attack AUC
Figure 12: Attack accuracy and AUC on CIFAR-10 against three models with three different architectures. We report the highest attack accuracy and attack AUC across all attacks. The prediction accuracy delta indicates the prediction accuracy gap compared to the undefended models, with negative numbers indicating a decrease in model accuracy.
A-A Experimental Results on Different Model Architectures
This section presents our additional experimental results on different model architectures, including DenseNet121 [28], VGG16 [29], and ViT [30] [^1] ViT contains layers that are incompatible with the DPSGD implementation in the Opacus library, so we omitted DPSGD from the experiments on ViT.. The results are illustrated in Figure 12. Consistent with our findings on ResNet18, Diffence effectively reduces both attack accuracy and attack AUC without impacting model accuracy. For example, Diffence reduces the attack AUC on average by 5.3% and the attack accuracy by 4.6% on average on the VGG16 model. On the ViT model, Diffence achieves reductions of 7.7% in attack AUC and 6.6% in attack accuracy on average. Additionally, it can be observed that combining Diffence with recent defenses such as SELENA and HAMP achieves the best privacy-utility trade-off.
(a) Attack Accuracy
(b) Attack AUC
Figure 13: Attack accuracy and AUC on three datasets against ResNet18, where Diffence employs a publicly released diffusion model trained on ImageNet.
Figure 14: Examples of original samples and their reconstructions on CIFAR-10, with reconstructions generated by a diffusion model pretrained on ImageNet.
TABLE VI: Comparison of Label-only MIAs against ResNet18 w/o and w/ Diffence.
Defenses Gap Attack Boundary Attack
Accuracy w/o Diffence (%) Accuracy w/ Diffence (%) AUC w/o Diffence (%) Accuracy w/o Diffence (%) AUC w/ Diffence (%) Accuracy w/ Diffence (%)
Undefended 58.5 58.5 70.8 69.5 70.9 69.3
SELENA 52.9 52.9 59.7 58.9 59.7 59.4
AdvReg 57.6 57.6 61.7 59.7 60.9 59.4
HAMP 58.6 58.6 69.5 66.8 69.5 67.3
RelaxLoss 58.2 58.2 67.6 66.3 67.7 66.4
DPSGD 54.6 54.6 57.3 56.5 57.7 57.1
MemGuard 58.5 58.5 70.8 69.5 70.9 69.3
A-B Impact of Diffence on label-only MIAs
In this section, we discuss the impact of Diffence on label-only MIAs [20, 21] and defenses. Both Diffence and label-only attacks rely on perturbations to the original images. Intuitively, Diffence modifies the input only after the target model $f$ has assigned a label to the noised inputs $x^{\prime}$ created by the label-only attack. This process alters the output confidence score without changing the predicted label $f(x^{\prime})$. Therefore, the target model with or without Diffence responds the same to label-only attack queries. Consequently, Diffence has no impact on label-only MIAs or defenses against them.
We used two types of label-only MIAs: the baseline gap attack [18] and the state-of-the-art boundary attack [21, 20] to test the impact of Diffence. As shown in Table VI, as expected, Diffence has no effect on the performance of label-only attacks and defenses against them. In all settings, the results with and without Diffence were nearly identical.
A-C Defending with Stronger Diffusion Models
TABLE VII: Performance of Diffence on CIFAR-10 when using a stronger diffusion model.
Defenses Test Accuracy (%) Attack Accuracy (%) Attack AUC (%)
Undefended / +Diffence 81.4 / 83.2 73.0 / 66.2 75.9 / 69.3
AdvReg / +Diffence 78.4 / 80.3 60.3 / 57.5 61.0 / 58.2
DPSGD / +Diffence 74.7 / 75.7 55.3 / 54.4 55.3 / 54.5
SELENA / +Diffence 79.2 / 79.8 60.4 / 57.2 60.1 / 57.0
RelaxLoss / +Diffence 81.2 / 82.1 65.1 / 61.5 69.5 / 63.7
HAMP / +Diffence 81.1 / 82.7 65.2 / 62.5 69.4 / 62.2
In our previous experiments, we assumed that the training set size of the defender’s diffusion model and the size of the classifier’s training set are the same. Furthermore, to ensure no decrease in model accuracy, we only select reconstructed samples that match the original sample labels.
Interestingly, however, we discovered that when a stronger diffusion model is employed in the defense—specifically, a model trained on a dataset larger than that of the classifier, directly averaging the predictions of all generated samples can enhance both the classifier’s privacy and its accuracy.
We utilized a publicly available diffusion model trained on 50,000 samples from CIFAR-10 to defend our classifiers, which were trained on 25,000 samples. The results, as shown in the Table VII, indicate that in all cases, the classifiers not only achieved enhanced privacy but also exhibited higher test accuracy. Given that diffusion model training does not require labeled data, it can be performed on a large, public, unlabeled dataset. We leave a more detailed discussion on how diffusion models can improve classifiers’ accuracy for future work.
A-D Comparison of Defense Training Overhead
In addition to the inference overhead discussed in Section V-A, we also report the overhead of training-phase defenses during the training stage. The results are shown in Table VIII.
TABLE VIII: Training overhead comparison of different defenses.
Defenses CIFAR-10 (h) CIFAR-100 (h) SVHN (h)
Undefended 0.6 0.6 0.4
SELENA 10.9 11.9 9.3
AdvReg 10.9 12.6 6.8
HAMP 0.8 0.8 0.5
RelaxLoss 0.5 0.6 0.4
DP-SGD 0.8 0.9 0.6
A-E Additional Experimental Details
For all attacks, we randomly selected 2000 members and 2000 non-members as target samples to perform MIAs. For NN-based attacks, we used the same attack model architecture and training configuration as in previous work [6], saving training checkpoints and using the attack model that achieved the highest attack AUC on the target samples.
For defenses, HAMP [6] uses an entropy threshold $\gamma$ and a regularizer parameter $\alpha$, while RelaxLoss [5] uses $\alpha$ to balance utility and privacy. However, automated parameter selection is not provided. Following their papers, we used grid search to train multiple models with different parameters and selected the models based on their claimed impact on utility. HAMP includes both training-phase regularization and testing-time output modification (retaining predicted labels while randomizing confidence scores). In our experiments, we included only the training-phase defense of HAMP because its testing-time output modification renders the confidence scores meaningless, as the generated confidence scores are entirely derived from randomly generated images. Adversarial regularization [1] uses the $\alpha$ parameter to balance model accuracy and privacy protection. We set $\alpha$ to 6 for CIFAR-10, CIFAR-100, and SVHN, and 4 for CelebA and UTKFace. For DP-SGD [3], we used PyTorch Opacus  [^2] https://github.com/pytorch/opacus to train the DP-SGD model, fixing a norm clipping bound of 1.2 and setting a noise multiplier of 0.1 for CIFAR-10, 0.05 for CIFAR-100 and SVHN, and 0.01 for CelebA and UTKFace. For SELENA [4], we followed the original paper’ settings with $K=25$ and $L=10$, where $K$ is the total number of teacher models and $L$ is the number of teacher models whose training sets do not contain a given member sample.